# Mask Image Watermarking

### v2

@@ D.3 Scalability to Different Watermark Bit Lengths
Figure 12: Watermark extraction performance of MaskMark-D and MaskMark-ED with different bits length. We also show the results of WAM for comparison.
@@ D.4 Enhancing Robustness against Adaptive Attacks via Fast Fine-tuning
Figure 13: The effect of VAE fine-tuning on the robustness of MaskMark. Fine-tuning the VAE enhances robustness against VAE attacks, with minimal impact on the original robustness performance.
@@ D.6 Visualization Results of Global and Local Watermark Embedding
Figure 15: Visualization results of global watermark embedding using MaskMark-D and local watermark embedding using MaskMark-ED. The residual image is acquired by $10\times|I_{wm}-I_{orig}|$ for observation, highlighting the residual more clearly. The same residual visualization strategy is applied in the following figures as well.
@@ D.7 Visualization Results of Localization
Figure 16: Visualization results of watermark localization using different methods. The inpainting results are obtained by applying stable-diffusion-2-inpainting [16] to the masked regions for content reconstruction.
Figure 17: Visualization results of watermark localization using different methods. The inpainting results are obtained by applying stable-diffusion-2-inpainting [16] to the masked regions for content reconstruction.
Figure 18: Visualization results of watermark localization using different methods. The outpainting results are obtained by applying stable-diffusion-2-inpainting [16] to the masked regions for content reconstruction.
Figure 19: Visualization results of watermark localization using different methods. The outpainting results are obtained by applying stable-diffusion-2-inpainting [16] to the masked regions for content reconstruction.

### v3

@@ D.3 Scalability to Different Watermark Bit Lengths
Figure 12: Watermark extraction performance of MaskWM-D and MaskWM-ED with different bits length. We also show the results of WAM for comparison.
@@ D.4 Enhancing Robustness against Adaptive Attacks via Fast Fine-tuning
Figure 13: The effect of VAE fine-tuning on the robustness of MaskWM. Fine-tuning the VAE enhances robustness against VAE attacks, with minimal impact on the original robustness performance.
@@ D.6 Visualization Results of Global and Local Watermark Embedding
Figure 15: Visualization results of global watermark embedding using MaskWM-D and local watermark embedding using MaskWM-ED. The residual image is acquired by $10\times|I_{wm}-I_{orig}|$ for observation, highlighting the residual more clearly. The same residual visualization strategy is applied in the following figures as well.
@@ D.7 Visualization Results of Localization
Figure 16: Visualization results of watermark localization using different methods. The inpainting results are obtained by applying stable-diffusion-2-inpainting (rombach2022high, ) to the masked regions for content reconstruction.
Figure 17: Visualization results of watermark localization using different methods. The inpainting results are obtained by applying stable-diffusion-2-inpainting (rombach2022high, ) to the masked regions for content reconstruction.
Figure 18: Visualization results of watermark localization using different methods. The outpainting results are obtained by applying stable-diffusion-2-inpainting (rombach2022high, ) to the masked regions for content reconstruction.
Figure 19: Visualization results of watermark localization using different methods. The outpainting results are obtained by applying stable-diffusion-2-inpainting (rombach2022high, ) to the masked regions for content reconstruction.